# ULPT: Prompt Tuning with Ultra-Low-Dimensional Optimization

### v1

@@ 4.1 Experimental Settings
MRQA Others
Method #Param NQ HQA SQA NewsQA Avg. WG Yelp SciTail PAWS Avg.
Fine-tuning 220M 75.1 77.5 81.1 65.2 74.7 61.9 96.7 95.8 94.1 87.1
Adapter 1.9M 74.2 77.6 81.4 65.6 74.7 59.2 96.9 94.5 94.3 86.2
BitFit 280K 70.7 75.5 77.7 64.1 72.0 57.2 94.7 94.7 92.0 84.7
LoRA 3.8M 72.4 62.3 72.5 56.9 66.0 58.2 97.1 94.7 94.0 86.0
[…]
ATTEMPT 232K 70.4 75.2 78.5 62.8 71.4 57.6 96.7 93.1 92.1 84.9
PT 76.8K 70.0 74.7 75.3 63.0 70.8 49.6 95.6 92.0 57.9 73.8
DPT(r=10) 9.0K 71.3 75.5 76.3 63.5 71.7 49.6 96.1 95.6 92.2 83.4
DPT (r=256)  [^2] K 71.4 76.0 77.6 64.2 72.3 49.6 96.3 95.2 55.8 74.2
[…]
4.2 Main Results
@@
Performance on GLUE and SuperGLUE. As shown in Table 1, our ULPT achieves similar or higher performance on GLUE and SuperGLUE benchmark datasets compared with previous methods, while maintaining remarkable parameter efficiency.
@@
Profoundly, the ultra-low-rank configuration of $r=2$ retains at least $97\%$ performance of vanilla prompt tuning (PT), achieving average accuracy points of $84.0$ on GLUE and $71.2$ on SuperGLUE with only 2% of the parameters. This highlights the capability of ULPT and its advantage in large-scale LLM customization.
@@
With a moderate rank of $r=64$, our ULPT outperforms that with $r=256$ and other state-of-the-art models, showing that our approach not only reduces the number of parameters but also alleviates the overfitting problem. Specifically, the DPT model Xiao et al. (2023) learns an up-projection matrix, resulting in lower performance and 7x more parameters when the rank is controlled; even with the best rank $r=10$ suggested by the original paper Xiao et al. (2023), DPT is inadmissible as it is worse than our ULPT (with $r=64$) in both parameter efficiency and performance.
@@
Our ULPT also has significant advantages in multi-task setups. A transfer learning method initializes a model by task mixtures and then adapts it to a specific task; therefore, it cannot save parameters. Example studies of transfer learning include SPoT Vu et al. (2022) and ATTEMPT Asai et al. (2022). Our ULPT approach outperforms them in terms of accuracy and parameter efficiency, while offering a simpler training pipeline. Multi-task learning, on the other hand, shares certain parameters across different tasks Karimi Mahabadi et al. (2021); Ivison and Peters (2022); Wang et al. (2023), and thus, the parameter efficiency is measured on a per-task basis. Despite this, our ULPT still outperforms multi-task prompt tuning methods in both accuracy and per-task parameter efficiency due to its ultra-low-dimensional nature.
Performance on MRQA and other NLP tasks. Table 2 presents the results on the MRQA dataset and four additional tasks in the “Others” category. Following the standard practice on these benchmarks Wang et al. (2023); Shi and Lipani (2024), we run ULPT three times with different seeds and report the mean and standard deviation.
Unlike GLUE and SuperGLUE performance, ULPT exhibits consistent improvement when the rank is higher. This is probably because these tasks are more challenging, which aligns with the observation that full-model fine-tuning outperforms parameter-efficient methods on these tasks. Nevertheless, ULPT achieves competitive performance (slightly worse than the best-performing DePT approach), while saving parameters by multiple folds.
@@ 4.3 In-Depth Analyses
Figure 3: Left: Training loss curves on SST2 comparing ULPT with and without learnable shift and scale embeddings across different rank configurations. Right: Evaluation accuracy curves on SST2. For clarity, we present the case $r=2$, where our ULPT is at a disadvantage. The trend for other configurations is similar.
Figure 4: Pairwise similarities of the learned shift (left) and scale (right) embeddings for various rank configurations on SST-2.
Ablation study on shift and scale embeddings. We conduct an ablation study on the learnable shift embedding $\bm{b}\in\mathbb{R}^{d}$ and scale embedding $\bm{s}\in\mathbb{R}^{d}$, using the SST-2 dataset [^2] Our preliminary experiments show that prompt tuning on SST-2, a smaller dataset, leads to faster convergence, making it suitable for validating the effectiveness of the ablated models. However, for the rest of the analysis, we use MNLI and Natural Questions, to better test the expressiveness of the ablated models and reduce the risk of overfitting observed in our main experiment. with the T5-base model as the testbed, where we set the token number to be $n=100$. The results are shown in Figure 3. As seen, the dotted lines correspond to removing both shift and scale embeddings; their training loss remains high, suggesting that naïvely freezing the projection matrix $\tilde{{\bm{P}}}$ hinders the optimization process and consequently lowers the model performance. Introducing a learnable shift embedding $\bm{b}$ provides a substantial improvement (dashed lines), particularly in the low-dimensional configuration of $r=2$. A learnable scale embedding scale $\bm{s}$ further improves the training process and performance (solid lines). The ablation study shows that, although shift and scale embeddings are additional $2d$-many parameters, they play an important role in ultra-low-dimensional prompt tuning.
To further investigate the behavior of these embeddings, we analyze the pairwise cosine similarities of the shift $\bm{s}$ and scale $\bm{b}$ embeddings under different rank configurations, visualized in Figure 4. We see that they exhibit interesting patterns: the learned shift embeddings have consistently high similarity scores with different rank configurations, indicating their primary role as an alignment mechanism after up-projection. By contrast, the scale embeddings show near-zero pairwise similarities, as they depend on the sampled (and frozen) random projection matrix $\tilde{\bm{P}}$.
@@
Figure 5: Results on MNLI and Natural Questions with the T5-base model. The number of prompt tokens for both ULPT and naïve prompt tuning varies from $10$ to $100$.
Figure 6: Results on MNLI and Natural Questions with controlled numbers of trainable parameters, comparing ULPT and prompt tuning across three T5 model sizes (small, base, and large).
Analysis of prompt lengths and dimensions. Recall that Table 1 has analyzed our ULPT performance with different ranks. We now vary the number of prompt tokens and plot the trend in Figure 5. We see that our ULPT exhibits a similar trend as vanilla prompting, where the performance increases with a longer prompt. With an appropriate rank configuration, our ULPT consistently outperforms vanilla prompt tuning under different lengths.
@@
Our low-rank ULPT provides a trade-off between the prompt length and dimension. We compare ULPT with vanilla prompt tuning when the learnable parameters are controlled. For our ULPT, we keep the prompt token number as $100$ and vary the rank from $2$ to $256$; for vanilla full-dimensional prompt tuning, we vary the token number from $2$ to $50$. This analysis is also conducted with three model sizes: T5-small, T5-base, and T5-large.
@@
Figure 6 illustrates the results, showing that our low-dimensional ULPT with more tokens (solid lines) always outperform vanilla full-dimensional prompt tuning with fewer tokens (dashed lines). The analysis suggests that, when the number of learnable parameters is controlled, a longer prompt with a lower dimension offers more flexibility due to the additional Transformer steps.
Table 3: Results on MNLI and Natural Questions for training either ${\bm{P}}$ or ${\bm{Z}}$ with T5-base. Numbers in the brackets refer to the rank $r$ given the controlled number of parameters.
Train? #Trainable Parameters
@@
MNLI 81.9 (2) 82.9 (16) 84.9 (64) 85.5 (256)
82.9 (2) 84.5 (8) 85.3 (33)
NQ $\textbf{67.2}_{0.2}$ (2) $\textbf{68.0}_{0.4}$ (16) $\textbf{70.7}_{0.3}$ (64) $\textbf{72.6}_{0.2}$ (256)
66.9 (2) 70.0 (8) 72.0 (33)
Comparison with an alternative method of tuning ${\bm{P}}$. The low-rank decomposition ${\bm{E}}={\bm{Z}}{\bm{P}}$ allows an alternative approach that freezes ${\bm{Z}}$ and tunes ${\bm{P}}$, which contrasts with our approach that freezes ${\bm{P}}$ and tunes ${\bm{Z}}$. The comparison is shown in Table 3. The alternative setup (tuning ${\bm{P}}$) can be viewed as learning an up-projection from a set of random but frozen low-dimensional vectors. However, a key drawback of making ${\bm{P}}$ trainable is the rapid growth in the number of parameters when the rank $r$ increases, since $d\gg n$ in most practical scenarios. To ensure a fair comparison, we control the number of parameters by varying the rank $r$ for both methods.
@@
4.4 Results on Decoder Models
In our main experiments, we use the encoder–decoder T5 model Raffel et al. (2020), following most previous work on prompt tuning Lester et al. (2021); Wang et al. (2023); Shi and Lipani (2024).
We extend the evaluation of ULPT to Bloomz Muennighoff et al. (2023), a decoder-only model with three difference sizes: 560M, 1.7B, and 3B, having hidden dimensions of 1024, 2048, and 2560 respectively. For evaluation diversity, we select three mid-sized tasks from each task group: SST-2, HotpotQA, and Winogrande, providing assessment across classification, multi-hop reasoning, and coreference reasoning. Since Bloomz models are larger than the T5 series, we train up to 30K steps with a batch size of $4$, while keeping other hyperparamters the same as our main experiment.
We consider comparing ULPT with prompt tuning under different parameter budgets for text generation. Specifically, we vary the rank of ULPT from $2$ to $256$ while fixing the length $n=100$. For full-dimensional prompt tuning, the token number is adjusted to match the parameter count.
Results in Table 4 show that ULPT consistently outperforms prompt tuning across all model sizes and tasks given a fixed parameter budget. These findings align with our earlier analysis (§4.3), confirming that ULPT can be applied to different model architectures.

### v2

@@ 3.3 Theoretical Analyses
Table 1: Performance on GLUE and SuperGLUE benchmarks based on the T5-base model. We report standard evaluation metrics, namely, Pearson correlation for STS-B, F1 for MultiRC, and accuracy for other classification tasks. We replicate prompt tuning (PT; Lester et al., 2021) and DPT Xiao et al. (2023) using their default configurations. Our replicated PT results slightly exceed those reported in previous studies. All other baseline results are directly sourced from Shi and Lipani (2024). The suggested rank for DPT is r=10 based on Xiao et al. (2023); we replicate DPT with r=64 for controlled comparison with our ULPT. tTransfer learning methods. mMulti-task learning methods, whose “#param/task” scores are calculated based on the GLUE benchmark.
4 Experiments
4.1 Experiments on Language Understanding Tasks
Datasets.
We evaluate ULPT across $21$ NLP tasks in our main experiment, following prior work Asai et al. (2022); Wang et al. (2023); Shi and Lipani (2024). Those tasks are grouped into 4 categories: (1) GLUE is a benchmark suite consisting of various language understanding tasks, such as MNLI Williams et al. (2018), QQP Wang et al. (2018), QNLI Demszky et al. (2018), SST-2 Socher et al. (2013), STS-B Cer et al. (2017), MRPC Dolan and Brockett (2005), RTE Giampiccolo et al. (2007) and CoLA Warstadt et al. (2019). (2) SuperGLUE extends GLUE by considering more challenging tasks with limited training data, consisting of MultiRC Khashabi et al. (2018), BoolQ Clark et al. (2019), WiC Pilehvar and Camacho-Collados (2019), WSC Levesque et al. (2012), and CB De Marneffe et al. (2019). (3) The MRQA 2019 Shared Tasks are a set of QA tasks to test LLM generation capabilities, consisting of Natural Questions Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), SearchQA Dunn et al. (2017), and NewsQA Trischler et al. (2017). (4) Other classification tasks beyond the above test suites are also considered, including WinoGrande Sakaguchi et al. (2021), Yelp-2 Zhang et al. (2015), SciTail Khot et al. (2018), and PAWS-Wiki Zhang et al. (2019). Further details on these datasets are provided in Table 4 in Appendix A.1.
Baselines.
We evaluate ULPT against full-model fine-tuning, serving as a strong but parameter-intensive baseline. Second, we include state-of-the-art parameter-efficient methods such as Adapter Houlsby et al. (2019), AdapterDrop Rücklé et al. (2021), BitFit Ben Zaken et al. (2022), HyperFormer Karimi Mahabadi et al. (2021), HyperDecoder Ivison and Peters (2022), LoRA Hu et al. (2022), and Ladder Side-Tuning (LST; Sung et al., 2022). Third, we compare ULPT with vanilla prompt tuning (PT) and its variants: DePT learns offsets to the input token embeddings while using a separate learning rate for the prompt embeddings Shi and Lipani (2024), and DPT is closely related to ULPT as it decomposes prompt embeddings into low-rank matrices Xiao et al. (2023), but it differs from ours by learning the up-projection. Finally, we compare ULPT with transfer or multi-task learning methods, including SPoT Vu et al. (2022), ATTEMPT Asai et al. (2022), and MPT Wang et al. (2023).
Implementation details.
We use the T5-base model with $d=768$. Consistent with prior work Shi and Lipani (2024); Xiao et al. (2023), we set the number of prompt tokens $n=100$ for the prompt embeddings ${\bm{Z}}\in{\mathbb{R}}^{n\times r}$. For the rank $r$, we evaluate three ultra-low configurations $r=2$, $16$, $64$ that achieve over 90% dimensionality compression, and a more expressive configuration of $r=256$. We initialize the frozen up-projection matrix $\tilde{{\bm{P}}}$ with a standard normal distribution and find that ULPT is robust to alternative random initializations such as uniform. In our analysis (Section 4.3), T5-small ($d=512$) and T5-large model ($d=1024$) are considered to evaluate the generality of ULPT across different model sizes and input space dimensions. Further training details are provided in Appendix A.2.
Llama 1B Llama 3B
Method #Param VRAM Runtime GSM8K MBPP Avg. #Param VRAM Runtime GSM8K MBPP Avg.
ICL (4-shot) - - - 34.3 21.1 27.7 - - - 62.5 23.9 43.2
LoRA (r=1) 106.5K 10.22 215.1 38.5 26.7 32.6 286.7K 18.44 615.8 62.9 32.1 47.5
LoRA (r=4) 426.0K 10.22 215.7 40.1 27.2 33.7 1.15M 18.45 618.5 63.4 34.3 48.9
LoRA (r=8) 852.0K 10.23 216.5 40.2 24.7 32.5 2.29M 18.46 621.1 62.2 37.8 50.0
VeRA (r=1) 41.0K 10.02 216.0 39.3 24.4 31.9 114.7K 17.97 620.9 65.5 35.5 50.5
VeRA (r=4) 41.1K 10.02 216.6 39.6 27.8 33.7 114.9K 17.97 621.5 65.0 34.4 49.7
VeRA (r=8) 41.2K 10.02 217.0 40.9 29.5 35.2 115.1K 17.97 621.6 65.7 33.9 49.8
FourierFT (n=128) 4.1K 10.53 278.1 35.8 21.5 28.7 7.2K 19.80 880.9 63.1 21.9 42.5
FourierFT (n=512) 16.4K 10.53 278.5 34.9 27.3 31.1 28.7K 19.80 881.5 66.6 35.3 51.0
FourierFT (n=1024) 32.8K 10.54 278.5 36.6 25.9 31.3 57.3K 19.80 882.6 65.5 35.4 50.5
IA3 147.5K 10.83 229.2 39.7 26.4 33.1 286.7K 19.15 636.1 63.7 36.7 50.2
PT 20.5K 9.79 203.6 40.2 24.7 32.5 30.7K 17.18 589.5 65.3 33.1 49.2
ULPT (r=2) 4.1K 9.78 203.8 39.7 26.1 32.9 6.2K 17.17 589.2 65.6 35.9 50.8
ULPT (r=64) 4.7K 9.78 204.2 42.4 28.7 35.6 6.8K 17.17 589.6 65.4 36.5 51.0
ULPT (r=256) 6.7K 9.78 204.6 41.4 26.3 33.9 8.7K 17.17 589.6 66.4 34.6 50.5
Table 2: Efficiency and performance of the Llama 3.2 models on GSM8K and MBPP. We report training parameters, VRAM usages (GB), runtime (seconds/1K steps), GSM8K accuracy, and MBPP pass@1; In-context learning (ICL) serves as the lower-bound baseline.
Performance on GLUE and SuperGLUE.
@@ 4.1 Experiments on Language Understanding Tasks
As shown in Table 1, ULPT achieves similar or higher performance on GLUE and SuperGLUE benchmark datasets compared with previous methods, while maintaining remarkable parameter efficiency.
@@
Profoundly, the extreme configuration of $r=2$ retains at least $97\%$ performance of vanilla prompt tuning (PT) while saving 98% of the parameters. This highlights the capability of ULPT and its advantage in large-scale LLM customization. To the best of our knowledge, we are the first to demonstrate effective prompt tuning in a two-dimensional parameter space.
@@
When increasing the rank to $r=64$, ULPT outperforms that with $r=256$ and other state-of-the-art models. This suggests that ULPT effectively mitigates overfitting while remaining highly expressive. Specifically, the DPT model Xiao et al. (2023) learns an up-projection matrix, underperforms ULPT despite using 7x more parameters at the same rank; even with the best setting $r=10$ suggested by the original paper Xiao et al. (2023), DPT is inferior to ULPT with $r=64$ in both performance and efficiency.
@@
ULPT also exhibits clear advantages in multi-task setups. A transfer learning method initializes a model by task mixtures and then adapts it to a specific task; therefore, it cannot save parameters. Previous studies of transfer learning include SPoT Vu et al. (2022) and ATTEMPT Asai et al. (2022). Our ULPT approach outperforms them in terms of accuracy and parameter efficiency, while offering a simpler training pipeline. Multi-task learning, on the other hand, shares certain parameters across different tasks Karimi Mahabadi et al. (2021); Ivison and Peters (2022); Wang et al. (2023), and the parameter efficiency is measured on a per-task basis. Despite this, our ULPT still outperforms multi-task prompt tuning methods in both accuracy and per-task parameter efficiency.
Performance on MRQA and other classification tasks.
Following prior work on prompt tuning Lester et al. (2021); Wang et al. (2023); Shi and Lipani (2024), we evaluate ULPT on MRQA 2019 shared tasks as well as other language understanding tasks. Detailed results are provided in Appendix B.2. Consistent with our findings on GLUE and SuperGLUE, these results confirm that ULPT maintains competitive performance across diverse tasks while utilizing significantly fewer training parameters.
4.2 Experiments on Reasoning Tasks
Datasets and Baselines.
Our previous experiments (Section 4.1) evaluate ULPT on four diverse task suites using the encoder–decoder T5 model, following prior prompt-tuning work Lester et al. (2021); Wang et al. (2023); Shi and Lipani (2024).
We further extend our experiments to two reasoning tasks, GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021), using decoder-only Llama 3.2 models Meta (2024) at 1B and 3B scales, which have dimensionalities of $2048$, $3072$.
We compare ULPT against in-context learning Brown et al. (2020), vanilla prompt tuning, IA3 Liu et al. (2022a), LoRA and its recent ultra parameter-efficient variants: VeRA Kopiczko et al. (2024), which approximates the LoRA update matrices via random projections, and FourierFT Gao et al. (2024), which compresses weight updates by leveraging random spectral entries in the frequency domain. Implementation and training details are provided in Appendix A.3.
Performance on GSM8K and MBPP.
As shown in Table 2, ULPT achieves the best trade-off between efficiency and performance. With only a few thousand trainable parameters, ULPT uses the least VRAM, achieves faster training runtime, and consistently delivers higher accuracy across both tasks and model scales. At $r=64$, ULPT outperforms all baselines, including LoRA, VeRA, FourierFT, and vanilla prompt tuning.
By contrast, IA3, LoRA and its variants require orders of magnitude more parameters to match or even underperform ULPT. Although FourierFT can match ULPT’s parameter scale at $n=128$, it suffers from significantly lower accuracy, slower training speed, and higher memory usage.
Finally, ULPT incurs no additional inference overhead. Since the learned prompt tokens are prepended once and cached during autoregressive generation, the decoding throughput remains virtually unchanged, detailed analysis can be found in Appendix B.4.
@@ 4.3 In-Depth Analyses
Figure 3: Results with controlled numbers of trainable parameters, suggesting that ULPT’s longer prompt with lower dimensions offers more expressivity.
@@
Figure 4: Performance when the number of prompt tokens for both ULPT and naïve PT varies from $10$ to $100$.
Dimension–length trade-off drives expressivity.
@@
ULPT enables a trade-off between prompt length and dimension under a fixed parameter budget. To investigate this, we compare ULPT with vanilla prompt tuning when the learnable parameters are controlled. For ULPT, we fix the prompt token number at $100$ and vary the rank from $2$ to $256$; for vanilla full-dimensional prompt tuning, we vary the token number from $2$ to $50$. This analysis is conducted with two large datasets MNLI and Natural Questions across three model sizes: T5-small, T5-base, and T5-large.
@@
Figure 3 illustrates the results, showing that our low-dimensional ULPT with more tokens (solid lines) always outperforms vanilla full-dimensional prompt tuning with fewer tokens (dashed lines). The analysis suggests that, when the number of learnable parameters is controlled, a longer prompt with a lower dimension offers more expressivity due to the additional Transformer steps. We provide additional results in Appendix B.3, which confirms that this also holds for decoder-only models.
Longer prompts improve performance.
Recall that Table 1 has analyzed our ULPT performance with different ranks. We now vary the number of prompt tokens from $10$ to $100$ and plot the trend in Figure 4. We see that our ULPT exhibits a similar trend as vanilla prompting: performance increases with a longer prompt. With an appropriate rank configuration, our ULPT consistently outperforms vanilla prompt tuning under different lengths.
Figure 5: Left: Training loss curves comparing ULPT with no alignment (dotted), with learnable shift only (dashed), and with both shift and scale (solid). Right: Evaluation accuracy curves for ULPT at $r=2$. Adding shift significantly improves optimization and accuracy, while adding scale yields further gains. Trends are consistent across ranks.
Figure 6: Left: Shift embeddings learned with different ranks are highly similar, suggesting a general alignment role. Right: Scale embeddings vary significantly, indicating their dependence on frozen random projections.
Ablation on shift and scale embeddings.
We conduct an ablation study on the learnable shift embedding $\bm{b}\in\mathbb{R}^{d}$ and scale embedding $\bm{s}\in\mathbb{R}^{d}$, using the SST-2 dataset with the T5-base model as the testbed, where we set the token number to be $n=100$. The results are shown in Figure 5. As seen, the dotted lines correspond to removing both shift and scale embeddings; their training loss remains high, suggesting that naïvely freezing the projection matrix $\tilde{{\bm{P}}}$ hinders the optimization process and consequently lowers the model performance. Introducing a learnable shift embedding $\bm{b}$ provides a substantial improvement (dashed lines), particularly in the low-dimensional configuration of $r=2$. A learnable scale embedding scale $\bm{s}$ further improves the training process and performance (solid lines). The ablation study shows that, although shift and scale embeddings are additional $2d$-many parameters, they play an important role in ultra-low-dimensional prompt tuning.
To further investigate the behavior of these embeddings, we analyze the pairwise cosine similarities of the shift $\bm{s}$ and scale $\bm{b}$ vectors under different rank configurations, visualized in Figure 6. Interestingly, the learned shift embeddings show consistently high similarity scores with different rank configurations, indicating their primary role as a stable alignment mechanism after up-projection. By contrast, the scale embeddings show near-zero similarity, as they depend on the sampled (and frozen) random projection matrix $\tilde{\bm{P}}$.
Comparison with an alternative method of tuning projection matrix.
Train? #Parameters
@@
MNLI NF 82.9 (2) 84.5 (8) 85.3 (33)
$\textbf{81.9}_{0.1}$ (2) $\textbf{83.0}_{0.2}$ (16) $\textbf{84.9}_{0.1}$ (64) $\textbf{85.4}_{0.1}$ (256)
NQ NF 66.9 (2) 70.0 (8) 72.0 (33)
$\textbf{67.2}_{0.2}$ (2) $\textbf{68.0}_{0.4}$ (16) $\textbf{70.7}_{0.3}$ (64) $\textbf{72.6}_{0.2}$ (256)
Table 3: Training either low-dimensional embeddings $\bm{Z}$} (ours) or the up-projection $\bm{P}$} with T5-base. Numbers in the brackets refer to the rank $r$ given the controlled number of parameters. “NF” refers to non-feasible.
We follow the setup in Section 4.3 to ablate which component to tune in the low-rank decomposition. The low-rank decomposition ${\bm{E}}={\bm{Z}}{\bm{P}}$ allows an alternative approach that freezes $\bm{Z}$} and tunes $\bm{P}$}, which contrasts with our approach that freezes $\bm{P}$} and tunes $\bm{Z}$}. The comparison is shown in Table 3. The alternative setup (tuning $\bm{P}$}) can be viewed as learning an up-projection from a set of random but frozen low-dimensional vectors. However, a key drawback of making $\bm{P}$} trainable is the rapid growth in the number of parameters when the rank $r$ increases, since $d\gg n$ in most practical scenarios. To ensure a fair comparison, we control the number of parameters by varying the rank $r$ for both methods.